OpenStudy (okdutchman7):

Please help. WILL MEDAL! Ricky and his six sons sat around a circular table. How many seating arrangements were possible? @mathmale @triciaal @sleepyjess @satellite73 @jim_thompson5910 @zepdrix @robtobey @aaronq @Zale101

OpenStudy (ageta):

wait im ronaldo lol

satellite73 (satellite73):

if i am not mistake, for a circle it is \((n-1)!\)

OpenStudy (okdutchman7):

May you please explain more in depth? @satellite73

satellite73 (satellite73):

if it was not a circle, but a line, and you were arranging \(n\) people (or whatever) there would be \(n!\) ways to do it by the counting principle n choices for the first seat, n-1 for the second, n-2 for the third and so on

jimthompson5910 (jim_thompson5910):

Since you can rotate the circle however you like, this means that it does not matter who you place at the top of the circle. Let's say we always place person 1 at the top like shown here http://mathworld.wolfram.com/CircularPermutation.html so we go from n people to n-1 people. So we have n-1 people to arrange making (n-1)! possible permutations.

satellite73 (satellite73):

in a circle you would count the arrangement 1,2,3,4,5,6 the same as 2,3,4,5,6,1

OpenStudy (okdutchman7):

I'm confused. :3

OpenStudy (okdutchman7):

@mathmale

jimthompson5910 (jim_thompson5910):

`Ricky and his six sons ` there are 1+6 = 7 people Let's say Ricky is placed at the top of the circle as shown in the link. There would be (n-1)! = (7-1)! = 6! = 6*5*4*3*2*1 = 720 ways to arrange the 6 sons left over you're probably overthinking things? @okdutchman7

OpenStudy (okdutchman7):

So there are 6 ways in total? @jim_thompson5910

OpenStudy (okdutchman7):

I would think that there would be more than that.

OpenStudy (okdutchman7):

Oh wait 720 ways.

OpenStudy (mathmale):

Here I admit to going more by gut than by facts ( ! ), but submit that we are looking for the total number of combinations (arrangements) describing how 6 people could be positioned around a table. Ricky sits down first. One of the 6 sons would sit next to him. (Assume we continue to seat the sons to the left of Ricky or the previous son.) After that, one of the 5 remaining sons would sit next to the previously seated son. And so on. Using this line of reasoning, 6! strikes me as correct. Ricky is always there, so we worry only about how the 6 sons are seated.

OpenStudy (okdutchman7):

Okay...

OpenStudy (mathmale):

That works out to be 720 different seating arrangements.

OpenStudy (okdutchman7):

Okay well thank you.

jimthompson5910 (jim_thompson5910):

@okdutchman7 look at the link http://mathworld.wolfram.com/CircularPermutation.html at the examples of arranging 3 people and arranging 4 people There are 2 ways to arrange 3 people around a circle since (3-1)! = 2! = 2 There are 6 ways to arrange 4 people around a circle because (4-1)! = 3! = 6 The exclamation mark means factorial http://www.purplemath.com/modules/factorial.htm